OpenStudy (anonymous):

Jonah was keeping track of how many cars drove by every hour. He realized that the number of cars were in a sequence that followed the pattern n +3. How many cars drove by on the 9th hour? Answer 27 6 12 3

OpenStudy (pfenn1):

If n = the number of hours then what is n+3 for n=9?

OpenStudy (anonymous):

idk..

OpenStudy (pfenn1):

The number of cars = n+3 where n= the number of hours (9)